# Supervised Fine Tuning on Curated Data is Reinforcement Learning (and can be improved)

### v1

@@ 3.1 Notation and Background
we overload the notation of $\pi_{\text{ref}}$ for simplicity, where the expectation is over the reference $\pi_{\mathrm{ref}}(\tau)=p(s_{0})\prod_{t=0}^{T-1}p(s_{t+1}|s_{t},a_{t})\pi_{\mathrm{ref}}(a_{t}|s_{t})$ and we weight these samples via the importance weight $\nicefrac{{p_{\pi}(\tau;\theta)}}{{\pi_{\mathrm{ref}}(\tau)}}$.
@@ 3.2 SFT lower bounds RL
$J(\theta)\geq\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}\Big{[}\mathbb{I}(S(\tau)>0)\log p(\tau;\theta)\Big{]}=\underbrace{\mathbb{E}_{\tau\in\mathcal{D}^{+}}\Big{[}\log p(\tau;\theta)\Big{]}}_{\mathcal{J}_{\text{SFT}}(\theta)},$ (5)
where $\mathcal{D}^{+}=\{\tau_{i}|\tau_{i}\sim\pi_{\mathrm{ref}}(\tau),\mathbb{I}(S(\tau_{i})>0)\}_{i=1}^{N}$ is a dataset of trajectories sampled from $\pi_{\textrm{ref}}$ and filtered by the binary reward indicator. And we start optimization of $J(\theta)$ by initializing to the reference distribution, setting $\pi_{\theta_{0}}=\pi_{\mathrm{ref}}(\tau)$. If $\pi_{\mathrm{ref}}$ is not available in closed form we approximate it by first learning it via standard behavior cloning on the full dataset $\mathcal{D}$. Thus we can interpret SFT on filtered data as maximizing a lower bound on the RL objective.
@@
$J(\theta)\geq\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}\Big{[}\sum_{i=0}^{C}\mathbb{I}(S(\tau)>c_{i})\log p(\tau;\theta)\Big{]}\propto\underbrace{\mathbb{E}_{\tau\in\mathcal{D}_{Q}^{+}}\Big{[}\log p(\tau;\theta)\Big{]}}_{\mathcal{J}_{\text{SFT(Q)}}(\theta)},$ (6)
where $\mathcal{D}_{Q}^{+}=\mathcal{D}_{c_{1}}^{+}\cup\dots\cup\mathcal{D}_{c_{C}}^{+}$ with $\mathcal{D}_{c}^{+}=\{\tau_{i}|\tau_{i}\sim\pi_{\mathrm{ref}}(\tau),\mathbb{I}(S(\tau_{i})>c_{i})\}_{i=1}^{N}$. And we have the relation $\forall i:c_{i}>c_{i-1}$ as well as $c_{0}=0$. Hence, we can interpret SFT over quality curated data as maximizing a lower bound on the RL objective, and we have $J(\theta)\geq\mathcal{J}_{\text{SFT(Q)}}(\theta)\geq\mathcal{J}_{\text{SFT}}(\theta)$, thus one should consider training with data sampled proportional to quality annotations when available.
@@ 3.3 Improving SFT from the perspective of RL
Given that the bound SFT optimizes becomes looser as we train, thus weakening the connection to RL. We show that by re-framing the RL problem we can introduce a tighter bound that only requires a minimal change to the maximum likelihood objective used in SFT and can achieve better results.
@@
$J(\theta)$ $\geq\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}\Big{[}\frac{q(\tau)}{\pi_{\mathrm{ref}}(\tau)}\mathbb{I}(S(\tau))\log p(\tau;\theta)\Big{]}=\underbrace{\mathbb{E}_{\tau\in\mathcal{D}^{+}}\Big{[}\frac{q(\tau)}{\pi_{\mathrm{ref}}(\tau)}\log p(\tau;\theta)\Big{]}}_{\mathcal{J}_{\text{iw-SFT}}},$ (8)
where $\mathcal{D}^{+}=\{\tau_{i}|\tau_{i}\sim\pi_{\mathrm{ref}}(\tau),\mathbb{I}(S(\tau_{i}))>0\}_{i=1}^{N}$. Here, the auxiliary distribution $q(\tau)$ is now used to importance weight the examples in the dataset. And an analogous bound for quality scored data $\mathcal{D}_{Q}^{+}=\mathcal{D}_{1}^{+}\cup\dots\cup\mathcal{D}_{C}^{+}$ as in Eq. (6) can be found:
$J(\theta)$ $\geq\mathcal{J}_{\text{iw-SFT(Q)}}=\mathbb{E}_{\tau\in\mathcal{D}_{Q}^{+}}\Big{[}\frac{q(\tau)}{\pi_{\mathrm{ref}}(\tau)}\log p(\tau;\theta)\Big{]}.$ (9)
@@ 3.5 A Toy Example to build Intuition for SFT on curated/filtered data
[…]
@@ 4.1 Training LLMs for reasoning via SFT on curated data
Model # ex. AIME 2024 MATH 500 GPQA Diamond
@@ Appendix C Hyperparameters for LLM experiments
The hyperparameters of our model are set the same as in [Muennighoff et al., 2025]: we use cosine annealing for the learning rate, decaying to zero over a maximum of 5 epochs and using a warm-up ratio of 0.05. The target learning rate for AdamW [Loshchilov et al., 2017] is $51e-$ $\beta_{1}=0.9,\beta_{2}=0.95$ and the batch size is $8$. Weight decay is set to $41e-$ 4. We also use $g(x)=0.1\times\mathrm{clip}(x,0.2,1.8)$ to smooth/clip the importance weights. Given that we do not have access to the model weights for Gemini Flash (which was used to generate the reasoning traces), or its original training data, we also have to rely on an approximation for the reference distribution $\pi_{\mathrm{ref}}$ in the importance weights. Ideally, we would pick a model that is “close” to Gemini-Flash as an approximation. We tested multiple models at the 32B parameter level, comparing their summed log-probabilities over examples from the dataset and find that they assign similar likelihood (albeit giving different top-k next token predictions) and settled simply using the model we start from as the reference model such that $q,\pi_{\mathrm{ref}}$ and $\pi_{\theta}$ are all close at the beginning of training.
@@ Appendix E Dataset curation for control tasks
$\mathcal{D}=\{(\tau^{1},r_{1}),\cdots,(\tau^{n},r_{n})\}$, $Q$, $\mathcal{D}_{S}=\emptyset$, percentage cutoffs=$[c_{1},\cdots,c_{N}]$
@@ Appendix F Additional details for training on D4RL control tasks
For all training runs we use a linear warm-up for the learning rate followed by cosine annealing [Loshchilov and Hutter, 2017] we train with Adam [Kingma and Ba, 2015]. We also use a three layer MLP with 256 as hidden dimension. We first train a BC policy for 10000 steps with learning rate $31e-$ 3 and batch size 32. This policy is used as initialization for $\pi_{\theta}$. For subsequent fine tuning, we use batch size 256 for all tasks. For SFT(Q), we use learning rate 4e-5, warmup steps 300, total steps 4500, percentile cutoffs [90,95, 98], apart from the hopper-medium-expert-v2 where we use learning rate 4e-4. Hyperparameters for iw-SFT(Q) can be found in 4, we note that we first divide

### v2

@@ 3.1 Notation and Background
we overload the notation of $\pi_{\text{ref}}$ for simplicity, where the expectation is over the reference $\pi_{\mathrm{ref}}(\tau)=p(s_{0})\prod_{t=0}^{T-1}p(s_{t+1}|s_{t},a_{t})\pi_{\mathrm{ref}}(a_{t}|s_{t})$ and weight samples via the importance weight $\nicefrac{{p_{\pi}(\tau;\theta)}}{{\pi_{\mathrm{ref}}(\tau)}}$.
@@ 3.2 SFT lower bounds RL
$J(\theta)\geq\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}\Big{[}\mathbb{I}(S(\tau)>0)\log p(\tau;\theta)\Big{]}=c_{\mathrm{ref}}\underbrace{\mathbb{E}_{\tau\in\mathcal{D}^{+}}\Big{[}\log p(\tau;\theta)\Big{]}}_{\mathcal{J}_{\text{SFT}}(\theta)},$ (5)
where $\mathcal{D}^{+}=\{\tau_{i}|\tau_{i}\sim\pi_{\mathrm{ref}}(\tau),\mathbb{I}(S(\tau_{i})>0)\}_{i=1}^{N}$ is a dataset of trajectories sampled from $\pi_{\textrm{ref}}$ and filtered by the binary reward indicator and $c_{\mathrm{ref}}=\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}[\mathbb{I}(S(\tau)>0)]$ is constant independent of $\theta$. And we start optimization of $J(\theta)$ by initializing to the reference distribution, setting $\pi_{\theta_{0}}=\pi_{\mathrm{ref}}(\tau)$. If $\pi_{\mathrm{ref}}$ is not available in closed form we approximate it by first learning it via standard behavior cloning on the full dataset $\mathcal{D}$. Thus we can interpret SFT on filtered data as maximizing a lower bound on the RL objective.
@@
$J(\theta)\geq\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}\Big{[}\sum_{i=0}^{C}\mathbb{I}(S(\tau)>c_{i})\log p(\tau;\theta)\Big{]}=c_{\mathrm{ref}}^{Q}\underbrace{\mathbb{E}_{\tau\in\mathcal{D}_{Q}^{+}}\Big{[}\log p(\tau;\theta)\Big{]}}_{\mathcal{J}_{\text{SFT(Q)}}(\theta)},$ (6)
where $\mathcal{D}_{Q}^{+}=\mathcal{D}_{c_{1}}^{+}\cup\dots\cup\mathcal{D}_{c_{C}}^{+}$ with $\mathcal{D}_{c}^{+}=\{\tau_{i}|\tau_{i}\sim\pi_{\mathrm{ref}}(\tau),\mathbb{I}(S(\tau_{i})>c_{i})\}_{i=1}^{N}$. Note $c_{\mathrm{ref}}^{Q}=\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}[\sum_{i=0}^{C}\mathbb{I}(S(\tau)>c_{i})]$. And we have the relation $\forall i:c_{i}>c_{i-1}$. Hence, we can interpret SFT over quality curated data as maximizing a lower bound on the RL objective.
@@ 3.3 Improving SFT from the perspective of RL
The bound SFT optimizes becomes looser as we train, thus weakening the connection to RL. We show that by re-framing the RL problem we can introduce a tighter bound that only requires a minimal change to the maximum likelihood objective used in SFT and can achieve better results.
@@
$J(\theta)$ $\geq\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}\Big{[}\frac{q(\tau)}{\pi_{\mathrm{ref}}(\tau)}\mathbb{I}(S(\tau))\log p(\tau;\theta)\Big{]}=c_{\mathrm{ref}}\underbrace{\mathbb{E}_{\tau\in\mathcal{D}^{+}}\Big{[}\frac{q(\tau)}{\pi_{\mathrm{ref}}(\tau)}\log p(\tau;\theta)\Big{]}}_{\mathcal{J}_{\text{iw-SFT}}},$ (8)
where $\mathcal{D}^{+}=\{\tau_{i}|\tau_{i}\sim\pi_{\mathrm{ref}}(\tau),\mathbb{I}(S(\tau_{i}))>0\}_{i=1}^{N}$ and $c_{\mathrm{ref}}=\mathbb{E}_{\pi_{\mathrm{ref}}(\tau)}[\mathbb{I}(S(\tau)>0)]$. Here, the auxiliary distribution $q(\tau)$ is now used to importance weight the examples in the dataset. And an analogous bound for quality scored data $\mathcal{D}_{Q}^{+}=\mathcal{D}_{1}^{+}\cup\dots\cup\mathcal{D}_{C}^{+}$ as in Eq. (6) can be found:
$J(\theta)$ $\geq c_{\mathrm{ref}}^{Q}\mathcal{J}_{\text{iw-SFT(Q)}}=\mathbb{E}_{\tau\in\mathcal{D}_{Q}^{+}}\Big{[}\frac{q(\tau)}{\pi_{\mathrm{ref}}(\tau)}\log p(\tau;\theta)\Big{]}.$ (9)
@@ 3.5 A Toy Example to build Intuition for SFT on curated/filtered data
Consider a simple bandit optimization problem in which an agent has a choice of two actions $\mathcal{A}=\{\text{pull-left},\text{pull-right}\}$. The episode terminates after picking a single action. $R(\tau)$ is either 0 or 1; the probability distribution of $R(\tau)$ is defined by:
@@
Thus the optimal strategy is to always pull the right arm. We assume that we can observe the behavior of the system only by watching (collecting data from) a reference policy $\pi_{\mathrm{ref}}(a)=\mathcal{U}\{\text{pull-left},\text{pull-right}\}$ that picks between the two arms uniformly at random.
@@ 4.1 Training LLMs for reasoning via SFT on curated data
Model
# ex.
AIME 2024
MATH 500
GPQA Diamond
@@ Appendix C Hyperparameters for LLM experiments
The hyperparameters of our model are set the same as in [Muennighoff et al., 2025]: we use cosine annealing for the learning rate, decaying to zero over a maximum of 5 epochs and using a warm-up ratio of 0.05. The target learning rate for AdamW [Loshchilov et al., 2017] is 1e51e-$5\beta_{1}=0.9,\beta_{2}=0.9$5 and the batch size is $8$. Weight decay is set to 1e41e-4. We also use $g(x)=0.1\times\mathrm{clip}(x,0.2,1.8)$ to smooth/clip the importance weights. Given that we do not have access to the model weights for Gemini Flash (which was used to generate the reasoning traces), or its original training data, we also have to rely on an approximation for the reference distribution $\pi_{\mathrm{ref}}$ in the importance weights. Ideally, we would pick a model that is “close” to Gemini-Flash as an approximation. We tested multiple models at the 32B parameter level, comparing their summed log-probabilities over examples from the dataset and find that they assign similar likelihood (albeit giving different top-k next token predictions) and settled simply using the model we start from as the reference model such that $q,\pi_{\mathrm{ref}}$ and $\pi_{\theta}$ are all close at the beginning of training.
@@ Appendix E Dataset curation for control tasks
[…]
@@ Appendix F Additional details for training on D4RL control tasks
For all training runs we use a linear warm-up for the learning rate followed by cosine annealing [Loshchilov and Hutter, 2017] we train with Adam [Kingma and Ba, 2015]. We also use a three layer MLP with 256 as hidden dimension. We first train a BC policy for 10000 steps with learning rate 1e31e-3 and batch size 32. This policy is used as initialization for $\pi_{\theta}$. For subsequent fine tuning, we use batch size 256 for all tasks. For SFT(Q), we use learning rate 4e-5, warmup steps 300, total steps 4500, percentile cutoffs [90,95, 98], apart from the hopper-medium-expert-v2 where we use learning rate 4e-4. Hyperparameters for iw-SFT(Q) can be found in 4, we note that we first divide